### v3

Abstract
Advances in transformer-based language models have highlighted the benefits of language-specific pre-training on high-quality corpora. In this context, German NLP stands to gain from updated architectures and modern datasets tailored to the linguistic characteristics of the German language. GeistBERT seeks to improve German language processing by incrementally training on a diverse corpus and optimizing model performance across various NLP tasks. It was pre-trained using fairseq with standard hyperparameters, initialized from GottBERT weights, and trained on a large-scale German corpus using Whole Word Masking (WWM). Based on the pre-trained model, we derived extended-input variants using Nyströmformer and Longformer architectures with support for sequences up to 8k tokens. While these long-context models were not evaluated on dedicated long-context benchmarks, they are included in our release. We assessed all models on NER (CoNLL 2003, GermEval 2014) and text classification (GermEval 2018 fine/coarse, 10kGNAD) using $F_{1}$ score and accuracy. The GeistBERT models achieved strong performance, leading all tasks among the base models and setting a new state-of-the-art (SOTA). Notably, the base models outperformed larger models in several tasks. To support the German NLP research community, we are releasing GeistBERT under the MIT license.
keywords:
Language Modeling, Machine Learning for NLP, Resources and Evaluation
1 Introduction
The advancement of neural language modeling (LM) in natural language processing (NLP) has been driven by the development of contextual pre-trained word representations, particularly through transformer-based architectures. Models like Bidirectional Encoder Representations from Transformers (BERT) [9] have significantly impacted the field by providing robust, generalized representations that can be fine-tuned for specific downstream tasks, enhancing performance across various NLP applications. While much of the early work focused on English and multilingual models, it has become clear that single-language models, particularly those trained on large, high-quality corpora, can outperform their multilingual counterparts when applied to their target language.
Building on this understanding, the German NLP community has seen the introduction of models like GottBERT [26], which leveraged the German portion of the OSCAR [21] corpus to create a high-performance RoBERTa-based [18] model tailored specifically for the German language. However, as the field evolves, so too must the approaches to model training. Recent developments in pre-training methodologies, such as Whole Word Masking (WWM) [7] and the availability of newer, more extensive corpora like OSCAR23 [15], OPUS [28], and mC4 [34], present opportunities to further refine and enhance German language models.
RoBERTa-based models like GottBERT are often limited to a 512-token input length, which can be insufficient for long documents and broader contextual understanding. To address this, we extend the input length to 8k tokens using Longformer [2] and Nyströmformer [32], which mitigate the computational constraints of full self-attention. We adapted the models using roberta2longformer [^1] https://github.com/LennartKeller/roberta2longformer. Longformer replaces full self-attention with local and task-specific global attention, reducing complexity while preserving contextual awareness. It remains compatible with RoBERTa, allowing seamless weight transfer from pre-trained models [18]. Nyströmformer approximates self-attention using the Nyström method for low-rank matrix decomposition, eliminating the need for specialized attention patterns. This improves scalability and flexibility, without the need for additional pre-training, making it a strong alternative for handling longer sequences.
To fully leverage these architectural improvements for German NLP, we introduce GeistBERT, a German Enhanced Incremental Semantic Tuning BERT model. GeistBERT builds on the foundation laid by the best checkpoints of filtered GottBERT models (i.e. ${}^{\mathrm{f}}\mathrm{GottBERT}$), incrementally pre-training on top of it using a combination of partly deduplicated modern German corpora, including OSCAR23, mC4 from CulturaX [20], Wikipedia and corpora from OPUS. By integrating WWM and harnessing the diverse, expansive nature of these contemporary corpora, GeistBERT aims to set a new benchmark for German language models, offering improved performance and broader applicability across a range of NLP tasks.
Our contributions are as follows:
  • We incrementally trained GeistBERT on top of ${}^{\mathrm{f}}\mathrm{GottBERT}$ using a combination of partly deduplicated modern German corpora (OSCAR23, OPUS, mC4), OpenLegal and Wikipedia.
  • We integrated Whole Word Masking (WWM) into the pre-training process to enhance the model’s ability to capture semantic relationships within the German language.
  • We provide GeistBERT as a base model as well as Longformer and Nyströmformer variants to the community, accessible under an open-source license for further usage.
GeistBERT represents a significant step forward in the development of German-specific transformer models, offering enhanced capabilities through modern training techniques and expansive, high-varying data. By extending the input length to 8k tokens through Longformer and Nyströmformer, our model is well-suited for handling long-text tasks, enabling more effective document-level language understanding and processing.
2 Related Work
The rise of transformer-based models like BERT [9] marked a major shift in NLP, enabling significant performance improvements. Originally introduced as an English model and later as a multilingual version (mBERT), BERT’s success led to monolingual adaptations tailored to specific languages. For German, models like GermanBERT [^2] https://www.deepset.ai/german-bert and dbmdz BERT [^3] https://huggingface.co/dbmdz/bert-base-german-uncased emerged, trained on datasets of 12GB–16GB, sourced from Wikipedia, news articles, and legal texts.
RoBERTa enhanced BERT by training on a larger 160GB corpus, optimizing the architecture, and removing next sentence prediction. This strategy was applied to other languages, resulting in models like CamemBERT [19] for French and RobBERT [8] for Dutch, highlighting the benefits of large, diverse training corpora.
In German NLP, GBERT and GELECTRA [5] built on this progress by training on 145GB of the OSCAR corpus [21] and additional sources, surpassing earlier German BERT models. These advancements underscored the impact of larger, well-curated datasets on model performance. GottBERT further extended this development as one of the first German RoBERTa models, trained on the German OSCAR corpus. Its results demonstrated the importance of data diversity but also noted that excessive data cleaning might reduce corpus variance and affect downstream performance. GeistBERT refines this lineage by improving data variance, optimizing pre-training strategies, and enhancing resource efficiency, making it a more robust model for German NLP.
3 Methodology
Training Data and Pre-training
Compared to $\mathrm{GottBERT}$, GeistBERT was trained on a substantially larger corpus, totaling approximately 1.3TB of text data. The training data, listed in Table 1, was used in a shuffled version to enhance diversity and robustness during pre-training. GeistBERT was pre-trained using the same byte-level BPE tokenizer as $\mathrm{GottBERT}$, following the GPT-2 design with a vocabulary size of 52k. While the tokenizer architecture mirrors GPT-2, the vocabulary itself was trained from scratch on German text. Fairseq [23] was employed to compute the binary format for pre-training.
Using fairseq, we pre-trained the GeistBERT model on a highly variant corpus consisting of 1.3TB plain text data on 8 NVIDIA A40 GPUs. The model was trained with the RoBERTa base architecture for 100k update steps using a batch size of 8k, initializing the weights with ${}^{\mathrm{f}}\mathrm{GottBERT}_{base}$. We largely adhered to RoBERTa’s default training configuration [18], including dynamic masking, optimizer settings, and fixed sequence lengths (512 tokens). A 10k iteration warmup was applied, gradually increasing the learning rate to a peak of 0.0007, followed by a polynomial decay to zero. The pre-training process required approximately 8.3 days of computation time. In order to generate the Nyströmformer and Longformer version, we applied a post-hoc transformation to convert the resulting checkpoint for long-context use.
Table 1: Overview of the datasets used for training. The table lists the individual corpora, their sizes in gigabytes, their data sources, and whether they were deduplicated. The final corpus aggregates all listed datasets, resulting in approximately 1.3 TB of training data.
Corpus Documents Size (GB) Data Source Deduplicated
mC4 & OSCAR23 6,064,736,930 1316.57 CulturaX Yes
ELRC-4244, ELRC-4240, ELRC-4258, ELRC-4217, ELRC-4189, ELRC-4171, ELRC-4149 14,919,003 2.34 OPUS Yes
ECB 1,732,472 0.29 OPUS No
EUbookshop 18,203,612 2.34 OPUS No
Europarl 2,234,583 0.36 OPUS No
EuroPat 19,387,517 3.52 OPUS No
OpenSubtitles 41,612,280 1.35 OPUS No
TildeMODEL 5,059,688 0.79 OPUS No
German Wikipedia 4,767,776 7.23 Wikipedia No
OpenLegalData 209,526 2.48 OpenLegal No
Final corpus 6,172,863,387 1337.28
Downstream Tasks
We fine-tuned pre-trained BERT models using Huggingface [31] scripts, optimizing batch size and learning rate via grid search. NER and classification (CLS) tasks were trained for up to 30 epochs, while NLI tasks ran for up to 10 epochs using Fairseq-adapted hyperparameters. Each task was executed 24 times with varied hyperparameters, selecting the best checkpoint based on the highest $F_{1}$ score (accuracy for NLI). Performance was evaluated analogously to Scheible et al. [26] and compared with results from that study.
NLI
We evaluated NLI on the German XNLI dataset [6], an extension of MultiNLI [30], with 122k training, 2490 development, and 5010 test examples per language. Performance was measured by accuracy.
Named Entity Recognition
NER evaluation used the German CoNLL 2003 [29] and GermEval 2014 [3] datasets. CoNLL 2003 includes three entity classes, while GermEval 2014 provides fine-grained labels and nested markables. Both were evaluated using the $F_{1}$ score, with GermEval using an adapted metric accounting for label and span equality.
Text Classification
We evaluated classification on GermEval 2018 [24] (German tweet sentiment analysis) and 10kGNAD [25] (German news categorization). GermEval 2018 followed Chan et al. [5] splits, while 10kGNAD used a predefined 90%-10% split with 10% for validation. Both tasks were assessed using the mean $F_{1}$ score.
4 Results
GeistBERT establishes a new state-of-the-art (SOTA) in German NLP, outperforming both base and large models in multiple tasks while maintaining efficiency and robustness. The results of all the downstream tasks are listed in Table 4.
NLI
GeistBERTbase achieves 82.67% accuracy, surpassing all other base models. Its Longformer and Nyströmformer adaptations remain highly competitive, performing on par with or exceeding previous best-performing models.
Table 2: All the results of the experiments are shown in percent. They are all based on the test set and the best score out of 24 runs (selection based on validation set). While NLI is measured by accuracy, all the other metrics are $F_{1}$ measures. Best score in bold and second underlined, for large and base models respectively. For the $\mathrm{GottBERT}$ models, we report results on both the unfiltered and filtered corpora (marked by f). Each setup was evaluated per epoch, reporting the best (based on validation) and last checkpoints, with the latter marked by   . Values for non-GeistBERT models are taken from Scheible et al. [26].
Model XNLI GermEval 2014 CoNLL 03 GermEval 2018 10kGNAD
coarse fine
[…]
Named Entity Recognition
GeistBERTlongformer leads among base models, achieving 86.71% (CoNLL 03) and 88.45% (GermEval 2014), outperforming even some large models. GeistBERT’s strong results in both datasets validate its ability to capture complex entity relationships and maintain data diversity.
Text Classification
GeistBERT and its variants dominate classification tasks, achieving 80.56% (GermEval 2018 coarse) and 90.89% (10kGNAD). Notably, GeistBERTnyströmformer surpasses base and many large models in fine-grained classification, showcasing its effectiveness in nuanced text understanding.
5 Discussion
5.1 Principal Findings
The continued pre-training on GottBERT in combination with a broader, partly deduplicated German corpus (OSCAR23, OPUS, mC4, Wikipedia, OpenLegal) as well as the integration of WWM evidently leads to substantial improvements on various language modeling tasks. Effectively, this makes GeistBERT a novel SOTA model, surpassing even larger models in multiple German NLP tasks.
Following GottBERT’s widespread adoption in NLP research [27, 4, 17], GeistBERT extends its usability in machine translation [33], NER [13, 12], and domain-specific applications. In contrast to the TPU-based training used for GottBERT, GeistBERT was pre-trained on GPUs, which allowed for more flexible data preprocessing, including sentence-aligned segmentation. This made it possible to preserve natural sentence structure during training, even when using fixed-length sequences. Nevertheless, hyperparameter tuning remains a crucial factor for achieving strong downstream performance [10]. WWM contributed to improved tokenization, aligning with previous findings [19, 5]. Certain Longformer-based tasks required multiple GPUs, even with gradient accumulation [22] steps to simulate larger batch sizes (48 and 64). In our experience, Longformer consumed significantly more VRAM compared to Nyströmformer and RoBERTa, which could be trained on a single RTX 3090 (24 GB VRAM) without gradient accumulation. Beyond memory consumption, inference speed and training time are important practical considerations. Table 7 provides a detailed comparison of training and evaluation time across GeistBERT models and tasks. As shown, Longformer consistently incurred higher computational costs—both during training and inference—and often required multiple GPUs, whereas Nyströmformer offered a more efficient trade-off between resource usage and performance. While deduplication was applied to specific subcorpora, the overall dataset was not filtered, meaning some redundant or lower-quality data may still be present. Prior work suggests that models benefit from increased corpus diversity [19], and GeistBERT’s use of many different corpora likely contributed to its robustness. Noteworthily, we chose to continue pre-training from $\mathrm{GottBERT}$ rather than training GeistBERT from scratch, as it is common practice with domain-specific adaptations [17, 16, 1, 14]. This decision allowed us to leverage pre-trained weights and rely on existing German-specific tokenization. Additionally, given the limited time available for this project, we prioritized the training and thorough evaluation of a continued pre-training approach as a more conservative and focused choice over training multiple models including one from scratch. El Boukkouri et al. [11] argue that training from scratch with a specialized vocabulary could offer more tailored embeddings. Their findings also indicate that continued pre-training on a domain-specific corpus yields comparable results in most cases. Although our strategy of relying on GottBERT’s pre-work led to strong performance across multiple downstream tasks, we acknowledge that a comparison to training from scratch on the same data and architecture would provide additional insight. We consider this a direction for future work. Another limitation of our study is the absence of long-context evaluation. This is primarily due to the lack of established benchmarks for long-context understanding in German. Creating such datasets is a complex and resource-intensive task, and is typically pursued as part of dedicated efforts within the NLP community. We consider this an important avenue for future work.
5.2 Limitations
Several limitations should be acknowledged in this study. First, while deduplication was applied, it was only performed on specific subcorpora, not the entire dataset, potentially leaving redundant or noisy data in other parts of the corpus.
Additionally, GeistBERT’s training data, though diverse, remains specific to the selected corpora (OSCAR23, OPUS, mC4, Wikipedia, OpenLegal). Its generalization to other datasets or domains remains uncertain, and performance on dialects and cultural nuances within German may be limited. Further fine-tuning could improve adaptability to regional language variations.
Due to efficiency constraints and limited computational resources, we did not train a large GeistBERT model. While our results demonstrate the strong performance of the base models, larger architectures could potentially achieve even better results.
5.3 Ethical Considerations
As with all large-scale NLP models, GeistBERT inherits biases from its training data, which could impact downstream applications. While deduplication reduces noise, it does not eliminate underlying biases. Additionally, privacy concerns arise from large web-based datasets, as models may retain sensitive information. Future research should explore privacy-preserving techniques to mitigate this risk.
Another consideration is the environmental cost of training large transformer models. While we optimized for efficiency, pre-training and evaluation remains computationally expensive, contributing to high energy consumption and carbon emissions.
As a RoBERTa-based model, GeistBERT does not generate free-form text but can still propagate biases present in its training data, potentially affecting downstream applications such as classification, sentiment analysis, or automated decision-making. Ensuring fairness, transparency, and responsible usage is essential, especially in sensitive domains such as legal, medical, or financial NLP.
6 Conclusion
In this work, we present GeistBERT, a German RoBERTa-based model trained on a diverse, partly deduplicated corpus, incorporating Whole Word Masking (WWM) and extended input capabilities. GeistBERT achieves state-of-the-art (SOTA) performance, outperforming even larger models in multiple tasks. Our findings highlight the importance of corpus diversity and advanced pre-training techniques in model effectiveness. All GeistBERT models are released on Huggingface, with the base model additionally in fairseq, under the MIT license, advancing German NLP.
Funding Declaration
This research did not receive any external funding.
Competing Interests
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix A Model Properties
The number of parameters in BERT-like models varies significantly depending on their architecture (see Table A). The base version of BERT has approximately 110 million parameters, while the large version expands to 340 million. RoBERTa, an optimized variant with improved pre-training, features 125 million parameters in its base model and 355 million in its large model.
Multilingual models like XLM-RoBERTa are designed for cross-lingual tasks, with the base version containing 270 million parameters and the large version 550 million. Electra, using a generator-discriminator framework, achieves competitive performance with fewer parameters: around 110 million in the base model and 335 million in the large model.
For German NLP, GeistBERT introduces several architectural variants optimized for efficiency and extended context. The standard GeistBERT model contains approximately 125 million parameters, matching the configuration of GottBERTbase. The GeistBERTnyströmformer variant integrates sparse Nyström attention to support longer input sequences and improve scalability, resulting in a slightly larger model with 131 million parameters. Meanwhile, GeistBERTlongformer uses sliding-window attention to enable extended context lengths and has around 153 million parameters. Other widely used German models, such as GBERT, GELECTRA, and dbmdz BERT, adopt architectures comparable to BERTbase ($\sim$110M) and BERTlarge ($\sim$345M), but differ primarily in their pretraining corpora and learning objectives, rather than being specialized for a specific domain.
Table 3: The size of the vocabulary and the size of the parameters are shown for the model types used in this study. This table does not show other design differences of the models. Values were extracted using Huggingface’s transformers library.
Model Vocab Size #Params
[…]
Appendix B Perplexity
During the model pre-training the perplexity of the model is computed based on a test set for each optimization cycle (see Figure 1). After an initial sharp decrease, perplexity briefly increased for several steps before gradually declining until the final step. We assume that, given more training time, it would have continued to decrease further.
Figure 1: Perplexity of the GeistBERT model. Values based on the validation of each optimization cycle during the training.
Appendix C Parameters
Like GottBERT, we conducted a grid search with the parameter space detailed in Table 4. The optimal hyperparameters for each task, selected based on validation performance, are presented in Table C. The evaluation times are summarized in Table 3. All tasks were processed using two Nvidia RTX 3090 GPUs, leveraging Huggingface’s Transformers library (v4.34.1).
Table 4: Hyperparameters used in the grid search of the downstream tasks.
Parameter Values
Learning Rate 5e-5, 2e-5, 1e-5, 7e-6, 5e-6, 1e-6
Batch Size 16, 32, 48, 64
Epochs 30
Table 5: Computation time in hours and minutes for the downstream tasks summing up to 517 hours and 24 minutes, which are approximately 21.6 days.
Task Computation Time
XNLI 47:52
GermEval 2014 235:36
CoNLL03 92:45
GermEval 2018 coarse 45:46
fine 43:25
10kGNAD 85:49
Table 6: Hyperparameters of the best downstream task models for each task and pre-trained model. This table extends the original $\mathrm{GottBERT}$ setup by including GeistBERT models. BS refers to batch size, and LR denotes the learning rate.
Model GermEval 2014 CoNLL 03 GermEval 2018 10kGNAD
coarse fine
BS LR BS LR BS LR BS LR BS LR
[…]
Table 7: Training (TT) and evaluation runtime (ET) in seconds of the best downstream task models for each task. #GPU denotes the number of devices required for training and GA the gradient accumulation steps. The evaluation could be run on a single GPU.
Task GeistBERT GeistBERTnyströmformer GeistBERTlongformer
#GPU GA TT ET #GPU GA TT ET #GPU GA TT ET
XNLI 1 1 47,684 20.0 1 1 83,100 23.9 1 1 33,404 41.2
CoNLL 03 1 1 807 2.2 1 1 891 2.7 2 4 14,611 30.7
GermEval 2018 fine 1 1 200 9.0 1 1 754 4.0 1 1 4437 7.0
coarse 1 1 196 3.6 1 1 979 3.8 1 1 4788 8.8
GermEval 2014 1 1 2,553 2.6 2 1 19908 66.0 2 1 22644 54.8
10kGNAD 1 1 1,219 3.9 1 1 1,552 4.5 2 4 11,159 16.6