### v4

# GeistBERT: Breathing Life into German NLP

Abstract
Advances in transformer-based language models have highlighted the benefits of language-specific pre-training on high-quality corpora. In this context, German NLP stands to gain from updated architectures and modern datasets tailored to the linguistic characteristics of the German language. GeistBERT seeks to improve German language processing by incrementally training on a diverse corpus and optimizing model performance across various NLP tasks. We pre-trained GeistBERT using fairseq, following the RoBERTa base configuration with Whole Word Masking (WWM), and initialized from $\mathrm{GottBERT}$ weights. The model was trained on a 1.3 TB German corpus with dynamic masking and a fixed sequence length of 512 tokens. For evaluation, we fine-tuned the model on standard downstream tasks, including NER (CoNLL 2003, GermEval 2014), text classification (GermEval 2018 coarse/fine, 10kGNAD), and NLI (German XNLI), using $F_{1}$ score and accuracy as evaluation metrics. GeistBERT achieved strong results across all tasks, leading among base models and setting a new state-of-the-art (SOTA) in GermEval 2018 fine text classification. It also outperformed several larger models, particularly in classification benchmarks. To support research in German NLP, we release GeistBERT under the MIT license.
1 Introduction
The advancement of neural language modeling (LM) in natural language processing (NLP) has been driven by the development of contextual pre-trained word representations, particularly through transformer-based architectures. Models like Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019) have significantly impacted the field by providing robust, generalized representations that can be fine-tuned for specific downstream tasks, enhancing performance across various NLP applications. While much of the early work focused on English and multilingual models, it has become clear that single-language models, particularly those trained on large, high-quality corpora, can outperform their multilingual counterparts when applied to their target language.
Building on this understanding, the German NLP community has seen the introduction of models like GottBERT (Scheible et al., 2024), which leveraged the German portion of the OSCAR (Ortiz Suárez et al., 2020) corpus to create a high-performance RoBERTa-based (Liu et al., 2019) model tailored specifically for the German language. However, as the field evolves, so too must the approaches to model training. Recent developments in pre-training methodologies, such as Whole Word Masking (WWM) (Cui et al., 2021) and the availability of newer, more extensive corpora like OSCAR23 (Jansen et al., 2022), OPUS (Tiedemann, 2012), and mC4 (Xue et al., 2021), present opportunities to further refine and enhance German language models.
To fully leverage these developments for German NLP, we introduce GeistBERT, a German Enhanced Incremental Semantic Tuning BERT model. GeistBERT builds on the foundation laid by the best checkpoint of the filtered GottBERT model (i.e. ${}^{\mathrm{f}}\mathrm{GottBERT}_{base}$), incrementally pre-training on top of it using a combination of partly deduplicated modern German corpora, including OSCAR23, mC4 from CulturaX (Nguyen et al., 2023), Wikipedia and corpora from OPUS. By integrating WWM and harnessing the diverse, expansive nature of these contemporary corpora, GeistBERT aims to set a new benchmark for German language models, offering improved performance and broader applicability across a range of NLP tasks.
Our contributions are as follows:
  • We incrementally trained GeistBERT on top of ${}^{\mathrm{f}}\mathrm{GottBERT}_{base}$ using a combination of partly deduplicated modern German corpora (OSCAR23, OPUS, mC4), OpenLegal and Wikipedia.
  • We integrated WWM into the pre-training process to enhance the model’s ability to capture semantic relationships within the German language.
  • We provide GeistBERT as base model to the community, accessible under an open-source license for further usage.
GeistBERT represents a significant step forward in the development of German-specific transformer models, offering enhanced capabilities through modern training techniques and high-varying data.
2 Related Work
The rise of transformer-based models like BERT (Devlin et al., 2019) marked a major shift in NLP, enabling significant performance improvements. Originally introduced as an English model and later as a multilingual version (mBERT), BERT’s success led to monolingual adaptations tailored to specific languages. For German, models like GermanBERT [^1] https://www.deepset.ai/german-bert and dbmdz BERT [^2] https://huggingface.co/dbmdz/bert-base-german-uncased emerged, trained on datasets of 12GB–16GB, sourced from Wikipedia, news articles, and legal texts.
RoBERTa enhanced BERT by training on a larger 160GB corpus, optimizing the architecture, and removing next sentence prediction. This strategy was applied to other languages, resulting in models like CamemBERT (Martin et al., 2020) for French and RobBERT (Delobelle et al., 2020) for Dutch, highlighting the benefits of large, diverse training corpora and the use of language-specific vocabularies.
In German NLP, GBERT and GELECTRA (Chan et al., 2020) built on this progress by training on 145GB of the OSCAR corpus (Ortiz Suárez et al., 2020) and additional sources, surpassing earlier German BERT models. These advancements underscored the impact of larger, well-curated datasets on model performance. GottBERT further extended this development as one of the first German RoBERTa models, trained on the German OSCAR corpus. Its results demonstrated the importance of data diversity but also noted that excessive data cleaning might reduce corpus variance and affect downstream performance. GeistBERT refines this lineage by increasing data variance, optimizing pre-training strategies, and achieving strong performance without increasing model size, making it a robust and accessible model for German NLP.
3 Methodology
Training Data and Pre-training
Compared to $\mathrm{GottBERT}$, GeistBERT was trained on a substantially larger corpus, totaling approximately 1.3TB of text data. Training data was shuffled to support uniform sampling and minimize order effects during pre-training. GeistBERT was pre-trained using the same byte-level BPE tokenizer as $\mathrm{GottBERT}$, following the GPT-2 design with a vocabulary size of 52k. While the tokenizer architecture mirrors GPT-2, the vocabulary itself was trained from scratch on German text. fairseq (Ott et al., 2019) was employed to compute the binary format for pre-training.
Table 1: Overview of datasets used for training. The table lists the individual corpora, their sizes in gigabytes, their data sources, and whether they were deduplicated. The final corpus aggregates all listed datasets, resulting in approximately 1.3 TB of training data.
Corpus Documents Size (GB) Data Source Deduplicated
mC4 & OSCAR23 6,064,736,930 1316.57 CulturaX Yes
ELRC-4244, ELRC-4240, ELRC-4258, ELRC-4217, ELRC-4189, ELRC-4171, ELRC-4149 14,919,003 2.34 OPUS Yes
ECB 1,732,472 0.29 OPUS No
EUbookshop 18,203,612 2.34 OPUS No
Europarl 2,234,583 0.36 OPUS No
EuroPat 19,387,517 3.52 OPUS No
OpenSubtitles 41,612,280 1.35 OPUS No
TildeMODEL 5,059,688 0.79 OPUS No
German Wikipedia 4,767,776 7.23 Wikipedia No
OpenLegalData 209,526 2.48 OpenLegal No
Final corpus 6,172,863,387 1337.28
Using fairseq, we pre-trained the GeistBERT model on a highly variant corpus consisting of 1.3TB plain text data on 8 NVIDIA A40 GPUs. The model was trained with the RoBERTa base architecture for 100k update steps using a batch size of 8k, initializing the weights with ${}^{\mathrm{f}}\mathrm{GottBERT}_{base}$. We largely adhered to RoBERTa’s default training configuration (Liu et al., 2019), including dynamic masking, optimizer settings, and fixed sequence lengths (512 tokens). A 10k iteration warmup was applied, gradually increasing the learning rate to a peak of 0.0007, followed by a polynomial decay to zero. The pre-training process required approximately 8.3 days of computation time.
Downstream Tasks
We fine-tuned pre-trained BERT models using Huggingface (Wolf et al., 2019) scripts, optimizing batch size and learning rate via grid search. NER and classification (CLS) tasks were trained for up to 30 epochs, while NLI tasks ran for up to 10 epochs using fairseq-adapted hyperparameters. Each task was executed 24 times with varied hyperparameters, selecting the best checkpoint based on the highest $F_{1}$ score (accuracy for NLI). Performance was evaluated analogously to Scheible et al. (2024) and compared with results from that study.
NLI
We evaluated NLI on the German XNLI dataset (Conneau et al., 2018), an extension of MultiNLI (Williams et al., 2018), with 122k training, 2490 development, and 5010 test examples per language. Performance was measured by accuracy.
Named Entity Recognition
NER evaluation used the German CoNLL 2003 (Tjong Kim Sang and De Meulder, 2003) and GermEval 2014 (Benikova et al., 2014) datasets. CoNLL 2003 includes four entity types, while GermEval 2014 provides fine-grained categories and supports nested annotations. Both were evaluated using the $F_{1}$ score, with GermEval using an adapted metric accounting for label and span equality.
Text Classification
We evaluated classification on GermEval 2018 (Risch et al., 2018) (German tweet sentiment analysis) and 10kGNAD (Schabus et al., 2017) (German news categorization). GermEval 2018 followed the data splits defined by Chan et al. (2020), while 10kGNAD used a predefined 90%-10% train–test split, with 10% of the training set further held out for validation. Both tasks were evaluated using the mean $F_{1}$ score.
4 Results
GeistBERT sets a new state-of-the-art among base models for German NLP, outperforming all comparable models and closely approaching large-scale model performance across tasks. It even achieves absolute SOTA in GermEval 2018 fine-grained classification (see Table 4).
NLI
[…]
Model XNLI GermEval 2014 CoNLL 03 GermEval 2018 10kGNAD
coarse fine
[…]
Named Entity Recognition
GeistBERT achieves strong $F_{1}$ scores on both CoNLL 2003 (86.17%) and GermEval 2014 (88.47%), outperforming all other base models in our evaluation. It also surpasses all large-scale $\mathrm{GottBERT}$variants on GermEval 2014 and comes remarkably close on CoNLL 2003, with only a 0.11% gap to the lowest-scoring large variant. While top-performing large models such as GBERTlarge (87.19% on CoNLL) and XLM-Rlarge (88.83% on GermEval) remain ahead, GeistBERT narrows the performance gap significantly, demonstrating robust entity representation capabilities despite its compact size.
Text Classification
GeistBERTbase achieves strong performance across all classification tasks, ranking first in GermEval 2018 fine-grained classification (66.42%), second in 10kGNAD (90.89%), and third in GermEval 2018 coarse (79.67%). It consistently outperforms all other base models and surpasses several large-scale models, particularly in the fine-grained setting. The results indicate that GeistBERT performs competitively across diverse classification benchmarks, despite being a base-sized model.
5 Discussion
The continued pre-training of GottBERT on a broader and partially deduplicated German corpus consisting of OSCAR23, OPUS, mC4, Wikipedia, and OpenLegal, together with the use of WWM, leads to clear improvements across multiple language modeling tasks. GeistBERT establishes a new state of the art among base models and achieves competitive results with larger models across multiple German NLP benchmarks.
In contrast to the TPU-based training used for GottBERT, GeistBERT was pre-trained on GPUs, which allowed for more flexible data preprocessing, including sentence-aligned segmentation. This made it possible to preserve natural sentence structure during training, even when using fixed-length sequences. Nevertheless, hyperparameter tuning remains a crucial factor for achieving strong downstream performance (Dodge et al., 2020). WWM contributed to improved tokenization, aligning with previous findings (Martin et al., 2020; Chan et al., 2020). While deduplication was applied to specific subcorpora, the overall dataset was not filtered, meaning some redundant or lower-quality data may still be present. Prior work suggests that models benefit from increased corpus diversity (Martin et al., 2020), and GeistBERT’s use of many different corpora likely contributed to its robustness. Additionally, vocabulary size plays a role in performance (Toraman et al., 2023), though ours remains well-optimized. Noteworthily, we chose to continue pre-training from $\mathrm{GottBERT}$ rather than training GeistBERT from scratch, as it is common practice with domain-specific adaptations (Lentzen et al., 2022; Lee et al., 2019; Arefeva and Egger, 2022; Gururangan et al., 2020). This allowed us to reuse German-specific tokenization and pre-trained weights, and to focus on training and evaluating a single, well-defined setup within time constraints. While training from scratch with a custom vocabulary may yield more tailored embeddings (El Boukkouri et al., 2022), prior work suggests that continued pre-training often achieves comparable results. A direct comparison between continued pre-training and training from scratch on the same architecture and corpus remains an interesting avenue for future work. Following the broad adoption of GottBERT in German NLP (Scherrmann, 2023; Bressem et al., 2024; Lentzen et al., 2022; Xu et al., 2021; Frei et al., 2022; Frei and Kramer, 2023), we hope GeistBERT will be similarly received and applied.
6 Conclusion
In this work, we introduced GeistBERT, a German RoBERTa-based language model trained on a diverse and partially deduplicated corpus, incorporating WWM to enhance pre-training. GeistBERT achieves SOTA performance among base models and even outperforms several larger models across multiple tasks. These results underscore the importance of corpus diversity and WWM in improving downstream performance. GeistBERT is released under the MIT license on Huggingface and fairseq.
Limitations
Several limitations should be acknowledged in this study. First, while deduplication was applied, it was only performed on specific subcorpora, not the entire dataset, potentially leaving redundant or noisy data in other parts of the corpus.
Additionally, GeistBERT’s training data, though diverse, remains specific to the selected corpora (OSCAR23, OPUS, mC4, Wikipedia, OpenLegal). Its generalization to other datasets or domains remains uncertain, and performance on dialects and cultural nuances within German may be limited. Further fine-tuning could improve adaptability to regional language variations.
Due to efficiency constraints and limited computational resources, we did not train a large version of GeistBERT, as pretraining based on GottBERT estimates would have required approximately 4.75 times more compute. While our results demonstrate the strong performance of the base model, larger architectures could potentially achieve even better results.
Ethical Considerations
Like all large-scale language models, GeistBERT may inherit biases from its training data, which can influence downstream tasks such as classification or decision-making. While deduplication reduces redundancy and noise, it does not remove deeper societal or representational biases. Furthermore, training on large web-based corpora raises privacy concerns, as models may inadvertently retain sensitive information. Responsible deployment is especially important in high-stakes domains like legal, medical, or financial NLP.
Despite optimizations for efficiency, pre-training and evaluating transformer models remain computationally demanding, contributing to energy use and carbon emissions. These environmental costs highlight the need for balancing model performance with sustainable development goals.
Appendix A Model Properties
Table A lists the vocabulary sizes and total parameter counts of all models included in our evaluation. While most German BERT-style base models, such as GBERTbase, dbmdzBERT, and GELECTRAbase, contain approximately 110 million parameters, GeistBERT and ${}^{\mathrm{f}}\mathrm{GottBERT}_{base}$ are slightly larger at around 126 million parameters due to their RoBERTa-based architecture and a larger vocabulary of 52,009 tokens.
Large-scale German models such as GBERTlarge, GELECTRAlarge, and ${}^{\mathrm{f}}\mathrm{GottBERT}_{large}$ contain between 335 and 357 million parameters. Among the multilingual models, XLM-RoBERTabase and XLM-RoBERTalarge are substantially larger, with 278 million and 560 million parameters respectively. The vocabulary sizes vary across models and are influenced by tokenizer design and pre-training data. GeistBERT uses the same tokenizer as GottBERT, which is based on byte-level BPE trained on German text.
Table 3: The size of the vocabulary and the size of the parameters are shown for the model types used in this study. This table does not show other design differences of the models. Values were extracted using Huggingface’s transformers library.
Model Vocab Size #Params
[…]
Appendix B Perplexity
During the model pre-training the perplexity of the model is computed based on a test set for each optimization cycle (see Figure 1). After an initial sharp decrease, perplexity briefly increased for several steps before gradually declining until the final step. We assume that, given more training time, it would have continued to decrease further.
Figure 1: Perplexity of the GeistBERT model. Values based on the validation of each optimization cycle during the training.
Appendix C Parameters
Like GottBERT, we conducted a grid search with the parameter space detailed in Table 4. The optimal hyperparameters for each task, selected based on validation performance, are presented in Table C. The evaluation times are summarized in Table 3. All tasks were processed using two Nvidia RTX 3090 GPUs, leveraging Huggingface’s Transformers library (v4.34.1).
Table 4: Hyperparameters used in the grid search of the downstream tasks.
Parameter Values
Learning Rate 5e-5, 2e-5, 1e-5, 7e-6, 5e-6, 1e-6
Batch Size 16, 32, 48, 64
Epochs 30
Table 5: Computation time in hours and minutes for the downstream tasks summing up to 517 hours and 24 minutes, which are approximately 21.6 days.
Task Computation Time
XNLI 47:52
GermEval 2014 235:36
CoNLL03 92:45
GermEval 2018 coarse 45:46
fine 43:25
10kGNAD 85:49
Table 6: Hyperparameters of the best downstream task models for each task and pre-trained model. This table extends the original $\mathrm{GottBERT}$ setup by including GeistBERT models. BS refers to batch size, and LR denotes the learning rate.
Model GermEval 2014 CoNLL 03 GermEval 2018 10kGNAD
coarse fine
BS LR BS LR BS LR BS LR BS LR
[…]